# Improved upper bounds on key invariants of Erdős-Rényi numerical semigroups

### v2

@@ 4. Proof of the Main Theorem
$\leq 1\cdot u(p)+\mathrm{Pr}[\lnot(D_{1}\land D_{2}\land D_{3}]\mathrm{E}[F|F\geq u(p)]$
$\leq\frac{4800}{p}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}+3\left(\frac{p}{24}\right)^{4}\cdot\frac{6\cdot 10^{9}}{p^{4}}$
$<\frac{4800}{p}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}+56000$
$<\frac{4800}{p}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}+\frac{70}{p}\left(\ln\left(\frac{24}{p}\right)\right)^{3}$
$<\frac{5000}{p}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}.$
We now turn to the embedding dimension. For any semigroup $\mathcal{S}$ with Frobenius number $f$, by definition the numbers $f+1,f+2,\dots,2f+1$ all belong to $\mathcal{S}$ and these numbers alone generate every number greater than $12f+$ 1. Thus no minimal generator is larger than $12f+$ 1, and also $f$ itself is not a generator, so a loose bound on the embedding dimension is $\operatorname{e}(\mathcal{S})\leq 2\operatorname{F}(\mathcal{S})$. That is, our random variables satisfy $Z\leq 2F$. We will apply this bound in the case that $F\geq u(p)$, and thus reuse the same calculation of the second term above.
[…]
@@
$\leq pu(p)+2\cdot 3\left(\frac{p}{24}\right)^{4}\cdot\frac{6\cdot 10^{9}}{p^{4}}$
$\leq 4800\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}+112000$
$\leq 4800\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}+2000\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}$
$<7000\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}.\qed$
@@ 5. Experiments, Conclusions, and Future Work
$\frac{K_{1}}{p}\leq\mathrm{E}[\operatorname{F}(\mathcal{S}(p))]\frac{K_{2}}{p}\left(\log_{2}\left(\frac{1}{p}\right)\right)^{3}$
@@
That is, it appears that both of our new upper bounds are too large by two log factors while the lower bounds of De Loera, O’Neill, and Wilburne are too small by a single log factor. Furthermore, as we show in Figures 1 and 2, the true constant factors appear to be small: between 2 and 3 for the Frobenius number and between 1 and 2 for the embedding dimension. Additionally, the same bounds appear to apply even when $p$ is relatively large.

### v3

@@ 4. Proof of the Main Theorem
$\leq 1\cdot u(p)+\mathrm{Pr}[\lnot D_{3}]\mathrm{E}[F|F\geq u(p)]$
$=O\left(\frac{1}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{3}\right)+o\left(p^{4}\right)O\left(\frac{1}{p^{4}}\right)=O\left(\frac{1}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{3}\right)$
We now turn to the expected embedding dimension. For any semigroup $\mathcal{S}$ with Frobenius number $f$, by definition the numbers $f+1,f+2,\dots,2f+1$ all belong to $\mathcal{S}$ and these numbers alone generate every number greater than $12f+$ 1. Thus no minimal generator is larger than $12f+$ 1, and also $f$ itself is not a generator, so a loose bound on the embedding dimension is $\operatorname{e}(\mathcal{S})\leq 2\operatorname{F}(\mathcal{S})$. That is, our random variables satisfy $Z\leq 2F$. We will apply this bound in the case that $F\geq u(p)$, and thus reuse the calculation of the second term above.
To handle the case where $F<u(p)$, let $Y$ denote the number of generators selected up to $2u(p)$. This is just a binomial random variable. If $F<u(p)$ then by the same argument as above, no minimal generator is greater than $2u(p)$. That is, $Z\leq Y$ in this case. We calculate
@@
$=O(pu(p))+o(1)=O\left(\left(\ln\left(\frac{1}{p}\right)\right)^{3}\right).\qed$
@@ 5. Experiments, Conclusions, and Future Work
$\frac{K_{1}}{p}\leq\mathrm{E}[\operatorname{F}(\mathcal{S}(p))]\leq\frac{K_{2}}{p}\left(\log_{2}\left(\frac{1}{p}\right)\right)^{3}$
@@
That is, it appears that both of our new upper bounds are too large by two log factors while the lower bounds of De Loera, O’Neill, and Wilburne are too small by a single log factor. Furthermore, as we show in Figures 1 and 2, the true constant factors appear to be quite small. Additionally, this asymptotic behavior appears to apply even when $p$ is relatively large.
@@
Acknowledgments: The authors thank Jesús de Loera, Christopher O’Neill, Adolfo Quiroz, and an anonymous referee of a previous draft of this manuscript for their helpful comments. Tristram Bogart was supported by internal research grant INV-2025-213-3438 from the Faculty of Sciences of the Universidad de los Andes.